Influenced and fixed and variable costs on Price and profit. In 19A the velasquez products company produced a machine that sold for $600$of which 450 represented cost of goods sold and $50 represented marketing and administrative expanses. The cost of goods sold was compromised of 40% material cost, 40% of labor cost and 20% of factory overhead. During 19A, 2000 machines were sold. During 19B, increases 20% in cost of materials and 25% in cost of labor are anticipated . The compny plans to rise the selling price to $675 per unit with resulting decrease of 40% in the number of units to be sold.

Required:
1) An income statement for 19B, indicating the new costs per unit. Assume that materials and labor cost will still equal 80% of the cost of good sold for 19B and marketing and administrative expenses are still $50 per unit.

2) A revised income statement for 19B, disregarding the 80% relationship of material and labor costs to cost to good sold. After the statement required in part (1) was prepared, it was ascertained that the 20% factory overhead in 19A consisted of $100,000 fixed and $80,000 variable expenses. The decrease in number of units to be sold in 19B does not influence the fixed cost.
Answer #1

1) Income Statement for 19B;

New Cost in 19B are as follows:

a) Material ( 20% more than in 19A) = 180+36(180*20%) = 216

b) Labour (25% more than in 19B) = 180+45(180*25%) = 225

c) Factory Overheads ( It is given that material and labour consist of 80% of cost of goods sold) = 110.25 {[(216 + 225)/80]*100}

Total COGS = 551.25 or 551

Total units sold in 19B = 1200 units [2000-(2000*40%)]

Income statement for 19B

Sales : (675 *1200) = $8,10,000

Less: COGS : (-551 * 1200) = $-6,61,200

Less: Mareketing and Admin: (-50 * 1200) = $-60,000

Profit: (74* 1200) = $88,800

2) Revised Income Statement

Revised Cost in 19B are as follows:

a) Material ( 20% more than in 19A) = 180+36(180*20%) = 216

b) Labour (25% more than in 19B) = 180+45(180*25%) = 225

c) Factory Overheads ( Total Factory overhead are $180,000):

(i)Fixed FOH = $100,000

(ii) Variable FOH = ($40 per unit * 1200) = $48,000

Note: As mentioned we have disregarded the 80% relationship of material and labour with COGS and hence variable overhead is taken as same $40 as in 19A.

Total units sold in 19B = 1200 units [2000-(2000*40%)]

Revised Income statement for 19B

Sales : (675 *1200) = $8,10,000

Less: COGS :   

(i)Material (216*1200)  = -$259,200

(ii)Labour (225*1200) = -$270,000

(iii)FOH (100000+48000) = -$148,000

Less: Mareketing and Admin: (-50 * 1200) = $-60,000

Profit: $132,800